Let Z be a standard normal random variable. Use the calculator provided, or this table, to determine the value of c. =P≤−0.68≤Zc0.7367 Carry your intermediate computations to at least four decimal places. Round your answer to two decimal places
Answer #1

Ane consider D(-0.68<?$C) = 0:7367. Zu NICI) 3 Plzsc)- PIZ=-0.68)= 0.7367 (C)- $(-0.68) -0.7367 $tz)zisiz) (C)= 0.7367- (068)​​​​